Correct Our environment is very sensitive to the amount of ozone in the upper atmosphere. The level of ozone normally found is 6.9 pa that the current ozone level is not at a normal level. The mean of 23 samples is 6.4 ppm with a sta A level of significance of 0.02 ill be used. Make the decision to reject or fail to reject the null hypothesis. ndard deviation of 1.1. Assume the population is normally distributea TablesKeypad Answer a Reject Null typothesis O Fall to Reject Null Hypothesis
Answer #1

The null and alternate hypothesis are:

H0: μ = 6.9

Ha: 11 FL 6.9

The test statistic value is given by: T_μ 6.4-6.9 =-2.18 1.1 Vn

Since this a two-tailed test, the p-value is given by:

2 × P(t <-2.18) 2 × P(t > 2.18) = 2 × 0.04024 = 0.08048

Since p-value > 0.02, so we do not have sufficient evidence to reject null hypothesis H0.

Thus, we say μ = 6.9.